OpenStudy (anonymous):

What is the value of s – 11.5 when s = –4.6?

OpenStudy (anonymous):

You plug in s into the equation of (s-11.5).

OpenStudy (anonymous):

So substitute -4.6 into (s-11.5)

OpenStudy (anonymous):

well i did -4.6-11.5= -16.1

OpenStudy (anonymous):

Then that's the answer

OpenStudy (anonymous):

thks

OpenStudy (texaschic101):

you got it :)

OpenStudy (anonymous):

No problem. I'd appreciate if you gave a medal :)

OpenStudy (anonymous):

could you help me with a few more?? if not thats ok

OpenStudy (anonymous):

Yea, I'm free

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

What're the problems?

OpenStudy (anonymous):

What value of z makes the equation true? 2.5 = –6.19 + z A. –8.24 B. –3.69 C. 4.31 D. 8.69

OpenStudy (anonymous):

This is like the last problem. So if you want to solve for z... what do you do first

OpenStudy (anonymous):

You need to balance your equation, and to solve for z, you want to have z on one side of the = sign, and everything else on the other side

OpenStudy (anonymous):

So, 2.5= -6.19 + z 2.5 + 6.19 = z

OpenStudy (anonymous):

What's 2.5 + 6.19